OpenStudy (anonymous):

I need to write this as sine cosine or tangent sin 9x cos x - cos 9x sin x

OpenStudy (anonymous):

Do I work backwards so Sin(a+b) where a = 9x and b = x?

OpenStudy (anonymous):

Sin10x

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Why not?

OpenStudy (anonymous):

cos it doesnt work like that without an equation

OpenStudy (anonymous):

sin 9x cos x - cos 9x sin x sin(a + b) = sin(a)cos(b) + cos(a)sin(b)

OpenStudy (anonymous):

So then its sin8x

OpenStudy (anonymous):

sin(a-b) = sin(a)cos(b ) - cos(a)sin(b) comparing to the above question, sin 9x cos x - cos 9x sin x =sin(9x-x)= =sin(8x)

OpenStudy (anonymous):

Ok